OpenStudy (anonymous):

[(2^x+1) x (8^-x)]=4

OpenStudy (zzr0ck3r):

8=2^3 so 2^(x+1) = 2^(-3x) = 4 log base 2 of both sides x+1-3x = 2 -2x=1 x=-1/2